You have been provided with the following information for Division X of a decentralized company:

Selling price $ 95
Variable cost per unit 62
Fixed cost per unit 29
Sales volume (units) 21,000
Capacity (units) 24,100


Division Y of the same company would like to purchase all of its units internally. Division Y needs 9,300 units each period and currently pays $89 per unit to an outside firm. What is the lowest price that Division X could accept from Division Y? (Assume that Division Y wants to use a sole supplier and will not purchase less than 9,300 from a supplier.)
Answer #1
Capacity (units) 24100
Less: Existing sales 21000
Excess capacity 3100
Existing sales lost 6200 =9300-3100
Variable cost of producing 9300 units 576600 =9300*62
Add: Contribution margin lost on existing sales 204600 =6200*(95-62)
Total minimum price to be charged 781200
Lowest price that Division X could accept from Division Y 84 =781200/9300